Find the maximun value of the function f(x)=x+x^2-x^3 for x>=0
1
Expert's answer
2019-09-23T06:10:54-0400

First of all, let's find derivative of our function:

"(x+x^2-x^3)'=1+2x-3x^2=0"

"D=4-4*(-3)=4^2, \\ x_1= \cfrac{-2-4}{-6}=1, x_2=\cfrac{-2+4}{-6}=-\cfrac{1}{3},"

"f'(1-\epsilon)>0, f'(1+\epsilon)<0
\implies x=1 -" local maximum, now let's prove that it's maximum value of the function when "x\in[0,+\infin): f(0)=0<f(1), f''(x)=2-6x \implies \forall x \in [1, +\infin): f''(x)<0," for this reason:

"f'(x)" is monotonically decreasing function  "\forall x \in [1, +\infin)" "\implies \forall x>1: f'(x)<0 \implies f(x) -" monotonically decreasing function. "f(1)=f_{max}(x)" for "x\in[0, +\infin)"